# biophysics, medical physics, biomechanics

### v1

@@ 3.3 Effects of varying spatial patterns of airway constriction
Both when the constrictions are clustered and when they are not, deposition typically increases in airways which are not directly proximal or distal to a constriction (figures 4c,g). This is because more of the inhaled gas is transported through these open paths once other paths are blocked, and flow through them is faster, so inertial impaction increases. When constrictions are clustered, increases in deposition in the central airways are generally larger than when constrictions are not clustered. This is evident in the comparison of figures 4(aiii) and 4(biii), and of figures 4(aiv) and 4(biv). Figure 4(aiv) shows positive, although relatively small, median increases in airway deposition in several generations of central airways. Figure 4(biv) shows almost exactly zero median change in the same generations when constrictions are not clustered. Variability in deposition in almost every generation of central airways is larger when constrictions are clustered. This highlights how clustered constrictions in the distal airways can cause more heterogeneous central-airways deposition.
@@
Figure 5: Data from simulations with six clusters of distal airway constriction (severity 0.9) in the upper lobes. In each simulation, six generation-12 airways were selected at random in the upper lobes, and all airways within a radius of $R_{C}$ of any of these were constricted, along with all of their descendants down to generation 15. Four realisations were simulated for each of $R_{C}=2.4\mathrm{cm}$, $R_{C}=3.2\mathrm{cm}$ and $R_{C}=4\mathrm{cm}$. (a) An example with $R_{C}=2.4\mathrm{cm}$, and (b) an example with $R_{C}=4\mathrm{cm}$, showing constricted airways (left) and scintigraphy (right). (c) Percentage change in deposition versus the unconstricted case, showing change in the upper lobes (left), and in the other lobes (right). These data are also separated into central conducting, distal conducting and acinar deposition. Mean $\pm$ standard deviation across the four realisations is plotted for each case. (d) Change in individual airway deposition for an example with $R_{C}=3.2\mathrm{cm}$.
@@
To examine the impact of increasing the size of clusters of constriction, we now apply $N_{C}=6$ clusters to the upper lobes, and vary the cluster radius, $R_{C}$ (figure 5). The process of generating clusters of constrictions is otherwise as described above. Simulated scintigraphy can be seen to clearly detect large clusters of constriction where many airways are blocked (figure 5b). Where clusters of constrictions are smaller and fewer airways are constricted, not all constricted regions are visible on scintigraphy (figure 5a). In figure 5(a), a patch of reduced deposition is clearly visible in the right lung, where two clusters are close together, but small clusters on the left lung, particularly those occurring in the centre of the image where lung depth is thickest, are not clear on scintigraphy. This highlights potential limitations of scintigraphy: it is more sensitive to severe blockages, and it may be more sensitive to constrictions that occur on the edges of the image field than those that occur in the centre.
@@ 4 Discussion
We have developed a model of inhaled particle deposition which predicts ventilation patterns based on the resistance of every airway in the lungs. We have tested the model in several lung geometries, validating total deposition results against the Multiple-Path Particle Dosimetry (MPPD) model [mppd_v3.04], showing good agreement even despite using different lung geometries. We investigated the impact of applying various patterns of distal airway constrictions, representative of features of CF disease. Strong changes in deposition can occur in the constricted airways themselves, but also, when changes to ventilation are taken into account, there can be significant impacts on deposition in unconstricted airways elsewhere in the lungs. Deposition in airways directly proximal or distal to constricted airways is reduced as flow rates through these paths are decreased. The severity of constriction at which this decrease becomes severe approximately coincides with a spike in the predicted LCI. We demonstrated that when distal airway constrictions are clustered together, their effects on deposition in the central airways can be heightened, with deposition becoming more heterogeneous. Central airways away from the constrictions received increased deposition, driven by enhanced inertial impaction as flow rates were increased through these open paths, since constricted airways received less flow and the flow rate into the trachea was assumed to be the same before and after applying constrictions. By contrast, the unconstricted distal conducting airways tended not to receive an increased dose since gravitational sedimentation, the dominant deposition mechanism in the small airways, is not increased by faster flow.
@@
Our model provides detailed predictions of deposition down to the single-airway scale, beyond the scope of current medical imaging techniques. The physical insights we have provided, relating changes in ventilation and deposition patterns, have the potential to improve estimates of regional dosing of an inhaled therapeutic once ventilation patterns have been inferred in a patient’s lungs from clinical tests of ventilation heterogeneity, such as MBW or ventilation MRI [horsley2009lung, smith2018patterns]. Dosing plans drawn up based on lung exposure calculated from MPPD and healthy volunteer studies may not be accurate for those with significant ventilation heterogeneity. Our data suggest there may be particular risks around higher concentrations of inhaled drug being delivered to central airways and lower doses to the most strongly affected lung regions.
The model has some limitations. Whilst the calculation of ventilation takes into account airway constriction, compliance of airways is not accounted for, so the impact of transient airway collapse and reopening during a breath is not described. This symptom of airway disease could impact deposition patterns, and so should be investigated in future studies. The model assumes that each acinus expands and contracts independently, so does not directly take into account how the motion of one acinus may modify its neighbours’ motion; this effect has been examined in previous ventilation models [roth2017comprehensive] and could be incorporated into future deposition models. Particle deposition in each airway is approximated using semi-empirical formulae, as has been done in many established whole-lung models [hofmann2011modelling]. CFD simulations may instead provide more accurate predictions of particle deposition in the central airways [oakes_aerosol_2017, kuprat2021efficient]. However, our approach keeps computational time low, meaning we have been able to simulate many different airway constriction scenarios, and make insights at the airway scale throughout the lungs. We have provided detailed physical insight into the effects of simulated patterns of airway constriction, representative of typical features of early CF disease. In future, patterns of airway constriction could be inferred directly from medical imaging, in order to move towards patient-specific predictions of the effects of disease. In order to isolate the impacts of airway constriction, we have not varied other parameters such as breathing rate or tidal volume. Faster, shallower breathing rates should be incorporated in future studies if modelling patients or patient-phenotypes where this is observed [hart2002changes]; this study has instead focused on features representative of mild disease, where breathing rates are unlikely to be significantly altered.
@@
Using a novel computational model that accounts for altered ventilation induced by airway constrictions, we have established new understanding of key impacts of airway disease on patterns of inhaled particle deposition. We have demonstrated how airway constrictions localised to the distal airways can affect deposition throughout the lungs. Spatial clustering of constrictions can increase the impact on central-airways deposition, making it more heterogeneous by reducing deposition in airways directly proximal to constrictions and increasing deposition elsewhere. These results have implications for understanding how ventilation defects and airway blockage in obstructive lung diseases may affect how inhaled therapeutics deposit. This is an important step towards developing realistic simulations of particle and drug deposition that can be used to accurately model exposure in disease states as well as healthy lungs.
@@
Code used to simulate ventilation and particle transport and deposition is available at xxxxxx. Code used to generate constriction patterns, and the network files used in all simulations presented in this study are available at xxxxxx. Code used to generate the scintigraphy plots, to calculate LCI, and to generate planar graphs from the 3D networks are available at xxxxxx. Generation of the lung networks used the same code as in [whitfield2020spectral].
@@
J.D.S. was supported by an EPSRC doctoral training award. A.H., C.A.W., and O.E.J. are supported by the NIHR Manchester Biomedical Research Centre (NIHR203308). The numerical algorithms developed extend on previous work by CAW funded by the UK Medical Research Council (MR/R024944/1)

### v3

@@ 3.3 Effects of varying spatial patterns of airway constriction
Both when the constrictions are clustered and when they are not, deposition typically increases in airways which are not directly proximal or distal to a constriction (figures 4aiii,biii). This is because more of the inhaled gas is transported through these open paths once other paths are blocked, and flow through them is faster, so inertial impaction increases. When constrictions are clustered, increases in deposition in the central airways are generally larger than when constrictions are not clustered. This is evident in the comparison of figures 4(aiii) and 4(biii), and of figures 4(aiv) and 4(biv). Figure 4(aiv) shows positive, although relatively small, median increases in airway deposition in several generations of central airways. Figure 4(biv) shows almost exactly zero median change in the same generations when constrictions are not clustered. Variability in deposition in almost every generation of central airways is larger when constrictions are clustered. This highlights how clustered constrictions in the distal airways can cause more heterogeneous central-airways deposition.
@@
Figure 5: Data from simulations with six clusters of distal airway constriction (severity 0.9) in the upper lobes. In each simulation, six generation-12 airways were selected at random in the upper lobes, and all airways within a radius of $R_{C}$ of any of these were constricted, along with all of their descendants down to generation 15. Four realisations were simulated for each of $R_{C}=2.4\mathrm{cm}$, $R_{C}=3.2\mathrm{cm}$ and $R_{C}=4\mathrm{cm}$. (a) An example with $R_{C}=2.4\mathrm{cm}$, and (b) an example with $R_{C}=4\mathrm{cm}$, showing constricted airways (left) and simulated scintigraphy (right). (c) Percentage change in deposition versus the unconstricted case, showing change in the upper lobes (left), and in the other lobes (right). These data are also separated into central conducting, distal conducting and acinar deposition. Mean $\pm$ standard deviation across the four realisations is plotted for each case. (d) Change in individual airway deposition for an example with $R_{C}=3.2\mathrm{cm}$.
@@
To examine the impact of increasing the size of clusters of constriction, we now apply $N_{C}=6$ clusters to the upper lobes, and vary the cluster radius, $R_{C}$ (figure 5). The process of generating clusters of constrictions is otherwise as described above. Since $N_{C}$ is kept fixed, increasing $R_{C}$ typically leads to a marked increased in the number of airways being constricted, so that the simulations with higher values of $R_{C}$ are representative of more severely diseased states compared to those with lower values of $R_{C}$. Simulated scintigraphy can be seen to clearly detect large clusters of constriction where many airways are blocked (figure 5b). Where clusters of constrictions are smaller and fewer airways are constricted, not all constricted regions are visible on simulated scintigraphy (figure 5a). In figure 5(a), a patch of reduced deposition is clearly visible in the right lung, where two clusters are close together, but small clusters on the left lung, particularly those occurring in the centre of the image where lung depth is thickest, are not clear on the simulated scintigraphy. This highlights potential limitations of scintigraphy: it is more sensitive to severe blockages, and it may be more sensitive to constrictions that occur on the edges of the image field than those that occur in the centre.
@@ 4 Discussion
We have developed a model of inhaled particle deposition which predicts ventilation patterns based on the resistance of every airway in the lungs. Existing computational models, such as the Multiple-Path Particle Dosimetry (MPPD) model [asgharian2001particle], have been built on assumptions that the lungs are healthy and that ventilation is uniform. Such models cannot predict the impacts of airway constriction and blockage on ventilation and particle deposition. Our model, therefore, provides a first step towards predicting particle deposition in patients with obstructive lung disease. The model outlined here provides a framework that can in future be extended to incorporate more clinically relevant physics, and to integrate clinical measurements of lung function and ventilation heterogeneity to move towards patient-specific predictions of drug deposition in obstructive lung diseases.
We have tested the model in several lung geometries, validating total deposition results in healthy lung simulations against MPPD [mppd_v3.04], showing good agreement even despite using different lung geometries. We investigated the impact of applying various patterns of distal airway constrictions, representative of features of CF disease. Strong changes in deposition can occur in the constricted airways themselves, but also, when changes to ventilation are taken into account, there can be significant impacts on deposition in unconstricted airways elsewhere in the lungs. Deposition in airways directly proximal or distal to constricted airways is reduced as flow rates through these paths are decreased. The severity of constriction at which this decrease becomes severe approximately coincides with a spike in the predicted LCI. We demonstrated that when distal airway constrictions are clustered together, their effects on deposition in the central airways can be heightened, with deposition becoming more heterogeneous. Central airways away from the constrictions received significantly increased deposition, driven by enhanced inertial impaction as flow rates were increased through these open paths, since constricted airways received less flow and the flow rate into the trachea was assumed to be the same before and after applying constrictions. By contrast, the unconstricted distal conducting airways tended not to receive an increased dose since gravitational sedimentation, the dominant deposition mechanism in the small airways, is not increased by faster flow. This disparity between the impact on central-airways and distal-airways deposition away from constrictions highlights a mechanism by which an increase in central-to-peripheral deposition ratio may be driven by the presence of localised severe distal airway constrictions, and our simulation predictions provide some quantification of this effect.
@@
Our model provides detailed predictions of deposition patterns throughout the central and distal airways, which have a significantly finer resolution than can be achieved with current medical imaging techniques. The physical insights we have provided, relating changes in ventilation and deposition patterns, have the potential to improve estimates of regional dosing of an inhaled therapeutic once ventilation patterns have been inferred in a patient’s lungs from clinical tests of ventilation heterogeneity, such as MBW or ventilation MRI [horsley2009lung, smith2018patterns]. Dosing plans drawn up based on lung exposure calculated from MPPD and healthy volunteer studies may not be accurate for those with significant ventilation heterogeneity. Our data suggest there may be particular risks around higher concentrations of inhaled drug being delivered to central airways and lower doses to the most strongly affected lung regions.
Some relevant physical effects are not included in the current model, but could be incorporated in future: for example, airway wall elasticity [venegas2005self], which could enable modelling of transient airway collapse; mechanical interdependence of the acini, which has been incorporated in previous ventilation models [roth2017comprehensive]; and extrathoracic deposition, which may be significant for large particles [darquenne_2012_aerosol]. Incorporating extrathoracic deposition could enable direct comparison with experimental data in patients; here, we have compared deposition results against MPPD simulations without extrathoracic deposition, which still provides good validation since MPPD compares well with experimental data [asgharian_2022_mixing, asgharian2006prediction]. To account for the effects of boundary-layer phenomena, plug-like flow, or the increase in resistance due to turbulent dissipation, particularly in the large airways, Poiseuille’s law could be replaced by an alternative resistance law [pedley1970prediction, reynolds1979modeling, lambert1982computational, ismail2013coupled]. The impacts of altered airway wall mechanics on particle deposition may be particularly relevant if extending the model to describe drug delivery in asthma [venegas2005self], and incorporating spatially varying elastic properties into the acinar model could enable modelling of heterogeneous emphysema in COPD [burrowes2014computational]. Whilst we have applied fixed patterns of airway constrictions in this study, incorporating a model for the formation of airway blockages, including modelling transient airway collapse and reopening, could be important for extending the model to capture a wider range of diseases and disease states in future.
We have provided detailed physical insight into the effects of simulated patterns of airway constriction, representative of typical features of CF. In order to isolate the impacts of airway constriction, we have not varied parameters such as breathing rate or tidal volume. Faster, shallower breathing should be incorporated if modelling patients or patient-phenotypes where this is observed [hart2002changes]; this study focused on features representative of mild disease, where breathing rates are unlikely to be significantly altered. We have demonstrated that the model can simulate deposition of particles of a wide range of sizes (figure 2e), but have focused on assessing the impacts of airway constriction on deposition of $4\mu\mathrm{m}$-diameter particles. Systematic investigation of the effects of varying particle size in lung models with airway constrictions is left for a future study. By assessing the impacts of breathing rates and particle size in diseased lungs, modelling results could then be used to suggest new clinical strategies for drug delivery.
The model presented here provides an advance on previous models, such as single-path [yeh1980models] or multiple-path [asgharian2001particle] models, by incorporating a physics-based ventilation model that accounts for the resistance of each individual airway. This means that it can predict the impacts of spatially heterogeneous patterns of airway constrictions or blockages on ventilation and particle deposition throughout the lungs. Models incorporating three-dimensional CFD simulations in the central airways provide more detailed predictions of deposition patterns, but the computational cost of CFD generally means only a small number of airways can be simulated [longest2012silico]. Experimental and numerical studies have shown that turbulence in the upper airways can enhance rates of particle deposition [darquenne2004aerosol], and CFD models may be able to capture these effects. Models that couple CFD in the large airways with reduced-order models for the distal airways can simulate deposition throughout the lungs on inhalation and exhalation [[, e.g., ]]oakes_aerosol_2017,kuprat2021efficient, but the whole-lung models of Oakes et al. [oakes_aerosol_2017, poorbahrami2021whole] or Kuprat et al. [kuprat2021efficient] assume uniform ventilation within the distal airways, so these models could not have responded to distal airway constrictions in a physics-based way, as our model does. Our reduced-order modelling approach lowers the computational time for our simulations compared to most CFD models, enabling us to run a larger number of simulations to investigate different patterns of airway constriction.
@@
Using a novel computational model that accounts for altered ventilation induced by airway constrictions, we have demonstrated some of the key impacts of airway disease on patterns of inhaled particle deposition. We have demonstrated how airway constrictions localised to the distal airways can affect deposition throughout the lungs. Spatial clustering of constrictions can increase the impact on central-airways deposition, making it more heterogeneous by reducing deposition in airways directly proximal to constrictions and increasing deposition elsewhere. These results have implications for understanding how ventilation defects and airway blockage in obstructive lung diseases may affect how inhaled therapeutics deposit. This is an important step towards developing realistic simulations of particle and drug deposition that can be used to accurately model exposure in disease states as well as healthy lungs.
@@
The code used to simulate ventilation and particle transport and deposition is available via xxxxxx. Code used to generate airway constriction patterns, the network files used in all simulations and the code used to generate simulated scintigraphy plots are available via xxxxxx. Generation of the lung networks used the same method as Whitfield et al. [whitfield2020spectral]; the code used to generate those networks was published previously here https://dx.doi.org/10.5281/zenodo.3709073, with the airway centre-line data from CT and computed networks here https://dx.doi.org/10.5281/zenodo.3709105.
@@
J.D.S. was supported by an EPSRC doctoral training award. C.A.W. is funded by, and A.H. and O.E.J. are supported by, the National Institute for Health and Care Research (NIHR) Manchester Biomedical Research Centre (BRC) (NIHR203308). A.H. and C.A.W. are supported by the ESPRC Network Plus “BIOREME” (EP/W000490/1). A.H. and C.A.W. also acknowledge funding from the Cystic Fibrosis Trust and Cystic Fibrosis Foundation through the Strategic Research Centre grant SRC025, which is supporting the continued development and translation of the research presented in this paper. The numerical algorithms developed extend on previous work by C.A.W. funded by the UK Medical Research Council (MR/R024944/1).